OpenStudy (anonymous):

Solve (y^-5)^-10y^10 y^-60 y^60 y^-150 y^150

jhonyy9 (jhonyy9):

so what is there (y^-5)^-10 *y^10 = ? correct ?

jhonyy9 (jhonyy9):

so in this case (y^-5)^-10 = y^50 *y^10 = y^(50+10) =y^60

jhonyy9 (jhonyy9):

do you understand it how ?